OpenStudy (anonymous):

How many 8 letter arrangements are possible using the letters of the word DECAGON if the letters A, E and O must occupy the third, fifth and sixth positions respectively and the letters remaining may be used more than once?

OpenStudy (anonymous):

4!

OpenStudy (anonymous):

sorry, that is wrong.

OpenStudy (cwrw238):

isn't it 5!

OpenStudy (anonymous):

nopee.

OpenStudy (anonymous):

Decagon has 7 letters. I counted as 8. Old age.

OpenStudy (cwrw238):

oh letters used more than once

OpenStudy (cwrw238):

yea - me too!

OpenStudy (anonymous):

\[ 4^5 \]

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

You have 5 slots left. You can have 4 choice for each slot. So you have \(4^5\) choices.

OpenStudy (anonymous):

that is also what i got, but the answer says 256 which is 4^4

OpenStudy (anonymous):

unfortunately there is five spaces.

OpenStudy (anonymous):

so 4^5, which is weird.

OpenStudy (anonymous):

as there are totally 8 places as three places are alredy fixed =1 *1*1=1 way so remaining 5 places can be may be used more than once in4^5 ways=1024 ways

OpenStudy (anonymous):

yes but the problem now is the answer says 256.

OpenStudy (anonymous):

The answer is wrong.